# TransNAS-TSAD: Harnessing Transformers for Multi-Objective Neural Architecture Search in Time Series Anomaly Detection [^1] This document is the results of the research project funded by the National Science Foundation.22footnotemark: 2

### 03 Dec 2023 v2

@@ Abstract
The surge in real-time data collection across various industries has underscored the need for advanced anomaly detection in both univariate and multivariate time series data. Traditional methods, while comprehensive, often struggle to capture the complex interdependencies in such data. This paper introduces TransNAS-TSAD, a novel framework that synergizes transformer architecture with neural architecture search (NAS), enhanced through NSGA-II algorithm optimization. This innovative approach effectively tackles the complexities of both univariate and multivariate time series, balancing computational efficiency with detection accuracy. Our evaluation reveals that TransNAS-TSAD surpasses conventional anomaly detection models, demonstrating marked improvements in diverse data scenarios. We also propose the Efficiency-Accuracy-Complexity Score (EACS) as a new metric for assessing model performance, emphasizing the crucial balance between accuracy and computational resources. TransNAS-TSAD sets a new benchmark in time series anomaly detection, offering a versatile, efficient solution for complex real-world applications. This research paves the way for future developments in the field, highlighting its potential in a wide range of industry applications.
@@
Time Series Anomaly Detection, Transformers in Machine Learning, Neural Architecture Search (NAS), NSGA-II Optimization, Deep Learning for Time Series Analysis, Adversarial Learning Paradigms
[…]
[1]organization=Department of Computer Science, addressline=University of Vermont,postcode=05405,city=Burlington, VT, country=U.S.A
@@ 1 Introduction
The prevalence of time series data in various sectors—from finance to healthcare, from infrastructure to manufacturing—highlights its critical role in modern analytics. Such data are often pivotal in uncovering underlying patterns and dependencies that are vital for strategic insights and anomaly detection, which can indicate significant shifts in system behaviors or emerging issues (Brophy et al., (2023); Li et al., (2023)).
@@
Traditional statistical methods have long been the mainstay for anomaly detection in time series data, trusted for their robust mathematical frameworks (Thudumu et al., (2020); Wang et al., (2011); Leadbetter, (1991)). Yet, the advent of Big Data—with its hallmark features of volume, velocity, and variety—has stretched these methods to their limits, revealing shortcomings through high false positive rates and missed detections (Ashabi et al., (2020); Al-Sai et al., (2019)).
@@
In contrast, deep learning has ushered in a paradigm shift, offering enhanced adaptability and performance in anomaly detection (Ismail Fawaz et al., (2019); Ma et al., (2021); Haq et al., (2023)). The transformer architecture, in particular, has emerged as a groundbreaking development due to its self-attention mechanism, enabling significant strides not only in natural language processing but also in time series analysis (Vaswani et al., (2017); Tuli et al., (2022); Kim et al., (2023)).
@@
However, the specific application of transformer models to time series anomaly detection necessitates further exploration to fine-tune their structures and parameters for optimal performance (Arslan et al., (2023)). Our initiative, TransNAS-TSAD, represents a significant shift from traditional transformer models, such as TranAD, to a more dynamic and adaptable framework. This new model integrates the transformer’s architectural strengths with systematic optimization through neural architecture search (NAS), emphasizing the multi-objective optimization strengths of the NSGA-II algorithm (Elsken et al., (2019); Deb et al., (2002)).
TransNAS-TSAD marks a significant evolution from TranAD’s static architecture, primarily through the integration of NSGA-II. This shift, grounded in (Lu et al., (2023)) study, leverages NSGA-II’s proven superiority in exploring complex search spaces more efficiently than recent algorithms. By adeptly balancing performance with computational efficiency, NSGA-II enhances our neural architecture design for anomaly detection, offering improved adaptability and efficiency. (Lu et al., (2020).
1.1 Research Context and Challenges
@@
Adapting anomaly detection models for time series data is a complex task, especially when dealing with multivariate series that present high-dimensional challenges and necessitate sophisticated model architectures (Wei, (2018); Liu et al., (2018)). The selective attention capabilities of the transformer model present a compelling solution, but only if the architectures are tailored to the specific characteristics of the dataset (Lian et al., (2019)).
@@
The marriage of NAS with multi-objective optimization algorithms like NSGA-II represents a strategic evolution in automated model design, offering a methodical path to discovering optimal architectures that consider multiple, often conflicting, objectives (Liu et al., (2021); Xue et al., (2023)).
1.2 Theoretical Foundations and Methodological Approach
Our methodology builds on the interplay between the theoretical principles of evolutionary algorithms and the empirical strengths of deep learning architectures. We aim to craft a robust NAS framework that can intelligently navigate and select transformer architectures with hyperparameters fine-tuned for anomaly detection in time series data (Wang et al., (2021); Ma et al., (2021)). The use of NSGA-II is central to this approach, providing a mechanism to balance the trade-offs between detection accuracy and model complexity (Wang et al., (2021)).
1.3 Guiding Philosophies and Contributions
Our research is guided by the following key principles:
  • 1.
    Adaptation to Time Series Data: We assert the need for specialized models that respect the unique temporal and dimensional characteristics of time series data.
  • 2.
    Maximizing Transformer Utility: We aim to exploit the full capabilities of the transformer architecture, particularly for temporal data interpretation.
  • 3.
    Strategic Optimization via NAS: We employ NAS, especially NSGA-II, as a systematic tool for refining architectures towards optimal model configurations.
Our contributions are as follows:
  1. 1.
    Novel Application of NAS: We leverage NAS to uniquely optimize transformer architectures for the domain of time series anomaly detection.
  2. 2.
    Architectural Parameter Exploration: We conduct extensive experimentation to identify transformer configurations that redefine design norms and improve detection performance.
  3. 3.
    Benchmarking Performance: We set new standards for performance, outperforming existing methods in a variety of datasets.
  4. 4.
    Model Design Efficiency: Our approach yields models that balance computational efficiency with analytical accuracy.
  5. 5.
    Architectural and Search Strategy Synergy: We seamlessly blend the structural advantages of transformers with the adaptive capabilities of NAS, resulting in models adept for time series anomaly detection.
@@ 2 Related Work
2.1 Anomaly Detection in Time Series Data
Anomaly detection in time series data has been a focal point of research due to its critical applications across various domains such as finance, healthcare, and industrial automation. Traditional approaches have primarily relied on statistical methods, which assume specific data distributions or apply distance-based metrics (Leadbetter, (1991); Wang et al., (2011)). However, these methods often struggle with high-dimensional, complex time series data inherent in modern applications (Hodge & Austin, (2004); Patcha & Park, (2007)).
With the advent of Big Data, researchers have sought more robust solutions capable of handling large volumes of high-velocity, high-variety data. Machine learning, and specifically deep learning, has emerged as a powerful approach to detect anomalies by learning representations of normal behavior and identifying deviations (Lee et al., (2022); Landauer et al., (2023); Al-amri et al., (2021)).
2.2 Transformers in Time Series Analysis
, The transformer architecture, introduced by Vaswani et al., (2017) has become a cornerstone for sequence modeling tasks, initially within the realm of natural language, processing and subsequently in time series analysis (Li et al., (2019); Tang & Matteson, (2021). Its self-attention mechanism allows it to capture temporal dependencies without the constraints of sequential data processing, distinguishing it from its predecessors like RNNs and LSTMs (Karita et al., (2019); Reza et al., (2022); Katrompas et al., (2022)).
In the domain of time series anomaly detection, transformers have shown promise due to their ability to model complex patterns over time (Tuli et al., (2022); Liu et al., (2021); Wang et al., (2022)). However, the research is still nascent, with considerable scope for optimization and adaptation to various time series characteristics (Luo et al., (2021); Wen et al., (2022)).
2.3 Neural Architecture Search (NAS)
NAS has emerged as a paradigm to automate the design of neural network architectures. Early NAS approaches were computationally intensive, employing reinforcement learning or evolutionary algorithms to explore the design space (Liu et al., (2021); Chen et al., (2019); Chu et al., (2020)). The development of efficient NAS methods has been a recent trend, focused on reducing the computational resources required for the search process (Jin et al., (2019); He et al., (2020)).
2.4 Multi-Objective Optimization in NAS
Multi-objective optimization in NAS aims to find a trade-off between competing objectives, such as model performance and computational cost. NSGA-II, an evolutionary algorithm, has been widely used for this purpose due to its efficiency in finding a diverse set of Pareto optimal solutions (Wang et al., (2020); Lu et al., (2019); Ying et al., (2020)). The application of multi-objective optimization in NAS for time series anomaly detection is still an under-explored area, offering a fertile ground for research (Borchert et al., (2022); Lu et al., (2021)).
2.5 Convergence of NAS and Transformers
The intersection of NAS and transformer models represents the cutting edge of research in the quest for optimal architectures for specific tasks. In time series anomaly detection, this convergence promises architectures that are both performant and efficient, addressing the unique challenges of temporal data (Wen et al., (2022)).
This work builds upon these foundations, aiming to address the identified gaps by leveraging the strengths of NAS and the transformer model to push the boundaries of anomaly detection in time series data.
@@ 3 Methodology
In this section, we delineate the comprehensive methodology employed in our study, designed to effectively detect and analyze anomalies in multivariate time series data using the TransNAS-TSAD framework. We commence with a clear definition of the problem, setting the stage for the subsequent methodological steps. This is followed by an in-depth discussion on data refinement techniques, ensuring that the data is appropriately pre-processed for our analysis. Next, we delve into the intricacies of our transformer architecture, highlighting its design and capabilities. A critical component of our approach is the neural architecture search (NAS) process, employing the NSGA-II algorithm for optimizing the transformer model, which we expound upon in detail. This is supplemented by an overview of the evolutionary process that iteratively refines our model architecture. We then transition to discussing advanced anomaly detection techniques, integrating adversarial elements into our model. The section culminates with a description of how we harness the full potential of TransNAS-TSAD for practical anomaly detection applications, illustrating the real-world implications of our research. Through this structured approach, we aim to provide a clear and logical flow of our methodology, ensuring that each step contributes coherently to achieving our research objectives.
@@ 3.1 Problem Definition
Given a sequence of multivariate time series data $S=\{x_{1},x_{2},\ldots,x_{T}\}$, where each point $x_{t}$ is timestamped $t$ and resides in an $m$-dimensional space ($x_{t}\in\mathbb{R}^{m}$), our aims are:
@@
3.1.1 Multi-Objective Transformer Model Optimization
At the heart of our methodology is the use of NSGA-II based neural architecture search (NAS) for optimizing the transformer architecture, focusing on:
  • 1.
    Maximizing Anomaly Detection Accuracy: We aim to enhance the F1 score for precise anomaly detection.
  • 2.
@@ 3.1.1 Multi-Objective Transformer Model Optimization
3.1.2 Data Processing and Model Training
Prior to modeling, the time series data undergoes a preprocessing phase where normalization is applied to ensure uniformity of scale across all series components. The normalization process is formulated as follows:
$x_{t}\leftarrow\frac{x_{t}-\min(S)}{\max(S)-\min(S)+\epsilon}$
where $\min(S)$ and $\max(S)$ denote the component-wise minimum and maximum values observed in the training time series $S$, respectively. The term $\epsilon$ is a small constant introduced to prevent division by zero, ensuring numerical stability.
@@ 3.1.2 Data Processing and Model Training
3.1.3 Anomaly Scoring and Thresholding
Our model is designed to assign an anomaly score to each context window $W_{t}$, as opposed to direct binary labels. This score is derived from the reconstruction error of the window, where $O_{t}$ denotes the reconstructed output corresponding to $W_{t}$. A dynamic threshold $D$ is established based on the distribution of the scores computed on the training data, facilitating the distinction between normal and anomalous windows. The precise mechanism for determining $D$ will be further elaborated upon in the subsequent sections, focusing on its adaptive nature and how it relates to the detection accuracy.
@@ 3.2 TransNAS-TSAD Transformer Architecture
In the TransNAS-TSAD framework, the transformer architecture is specially adapted for time series anomaly detection. This adaptation involves a unique configuration of the standard transformer model to align with the specific characteristics of time series data.
Data Augmentation Mechanisms Within the Model: TransNAS-TSAD integrates specific data augmentation techniques within the model architecture. These techniques, which can be fine-tuned during training, include:
  • 1.
@@
Encoder Operations: The encoder in TransNAS-TSAD is designed with the following operations:
  • 1.
@@
    $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{\text{model}}}}\right)V$ (1)
@@
    $\text{MultiHeadAtt}(Q,K,V)=\text{Concat}(H_{1},H_{2},\ldots,H_{h})$ (2)
@@
Decoder Operations: The decoder mirrors the encoder’s complexity and includes additional transformations that integrate encoder memory, apply multi-head self-attention, and utilize feedforward neural networks, culminating in a final layer for accurate reconstruction. TransNAS-TSAD can utilize either a dual sequential decoder setup or an iterative approach, depending on the optimized configuration for maximum anomaly detection accuracy.

### 04 Mar 2024 v3

@@ Abstract
The surge in real-time data collection across various industries has underscored the need for advanced anomaly detection in both univariate and multivariate time series data. This paper introduces TransNAS-TSAD, a framework that synergizes the transformer architecture with neural architecture search (NAS), enhanced through NSGA-II algorithm optimization. This approach effectively tackles the complexities of time series data, balancing computational efficiency with detection accuracy. Our evaluation reveals that TransNAS-TSAD surpasses conventional anomaly detection models due to its tailored architectural adaptability and the efficient exploration of complex search spaces, leading to marked improvements in diverse data scenarios. We also introduce the Efficiency-Accuracy-Complexity Score (EACS) as a new metric for assessing model performance, emphasizing the balance between accuracy and computational resources. TransNAS-TSAD sets a new benchmark in time series anomaly detection, offering a versatile, efficient solution for complex real-world applications. This research highlights the TransNAS-TSAD potential across a wide range of industry applications and paves the way for future developments in the field.
@@
Neural Architecture Search, Time Series Anomaly Detection, Transformers, Multi-Objective Optimization
 [^1] The source code, data, and/or other artifacts have been made available at https://github.com/ejokhan/TransNAS_TSAD.git.
@@ 1 Introduction
The ubiquity of time series data across various sectors, ranging from finance [10] and healthcare [62] to infrastructure [11] and manufacturing [31], underscores its pivotal role in modern analytics. These type of data are instrumental in identifying the patterns, dependencies, and anomalies indicative of significant shifts in system behaviors or the emergence of critical issues [12, 37].
@@
Traditionally, statistical methods have been the cornerstone for anomaly detection in time series data, and are renowned for their robust mathematical frameworks [60, 67, 49]. However, the advent of Big Data, characterized by its significant volume, velocity, and variety, has revealed the limitations of these traditional methods, including high false positive rates and missed detections [3, 6].
@@
In response, deep learning has introduced a paradigm shift, offering enhanced adaptability and performance in anomaly detection tasks [27, 52, 26]. Notably, the transformer architecture, with its self-attention mechanism, has emerged as a groundbreaking development, making significant strides not only in natural language processing but also in time series analysis [63, 59, 33].
@@
Despite the advantages offered by transformer models, their application to time series anomaly detection necessitates a more dynamic and adaptable framework for fine-tuning their structures and parameters for optimal performance [7]. Our model, TransNAS-TSAD, combines the architectural strengths of transformers (e.g., TranAD) with the optimization strategy of a neural architecture search (NAS) to systematically explore a huge number of architectural configurations. The latter leverages the multi-objective optimization capabilities of the NSGA-II algorithm [22, 21], enabling an efficient exploration of complex search spaces [45, 47]. This represents a substantial evolution of the TranAD’s static architecture, facilitating a balance between performance and computational efficiency.
@@
Adapting anomaly detection models to time series data is inherently complex, especially when addressing multivariate series that present high-dimensional challenges and necessitate sophisticated model architectures [68, 44]. The selective attention capabilities of the transformer model offer a compelling solution, provided the architectures are meticulously tailored to the specific research questions and dataset characteristics [46].
@@
The convergence of NAS with multi-objective optimization algorithms like NSGA-II marks a strategic evolution in automated model design. This approach provides a methodical pathway to discovering optimal architectures that consider multiple, often conflicting, objectives, thereby enhancing the adaptability and effectiveness of anomaly detection models [36, 70].
Our research is grounded on three fundamental principles designed to tackle the inherent challenges of time series anomaly detection. First, we emphasize the necessity of developing specialized models tailored to the unique temporal resolution and dimensional characteristics of time series data, ensuring they are finely attuned to the intricacies of temporal patterns. Second, we aim to maximize the utility of transformer architectures, exploiting their advanced capabilities for deep and nuanced interpretation of temporal data, which is critical for identifying subtle anomalies. Lastly, strategic optimization through Neural Architecture Search (NAS), particularly leveraging the NSGA-II algorithm, forms the cornerstone of our methodology. This approach allows us to systematically refine and perfect our models, ensuring they are not only highly effective but also optimized for the specific demands of the datasets they analyze. Together, these principles guide our pursuit of creating models that are both innovative and highly adapted to the complex domain of time series anomaly detection. Our contributions are delineated as follows:
  • We introduce a novel approach that applies NAS with NSGA-II optimization to specifically fine-tune transformer architectures for time series anomaly detection, showcasing NAS’s adaptability in specialized domains.
  • Our experimentation identifies key architectural configurations that elevate detection performance, highlighting the importance of architectural fine-tuning in enhancing model efficacy.
  • We demonstrate substantial performance improvements across diverse datasets, proving our optimized models’ superiority over existing methods and illustrating the benefits of multi-objective optimization for anomaly detection.
  • We strike an optimal balance between computational efficiency and analytical accuracy, presenting models that excel in real-time anomaly detection and are feasible for practical deployment, emphasizing efficiency in model design.
  • The synergy between transformers’ structural strengths and NAS’s adaptive optimization creates uniquely effective models for time series anomaly detection, marking a significant advancement by integrating architectural innovation with strategic optimization.
@@ 2 Related Work
The quest to detect and understand anomalies within time series data spans various domains including, but not limited to, finance [10], healthcare [62], infrastructure [11], and manufacturing [31]. This widespread interest underscores the fundamental role of anomaly detection in predictive maintenance, fraud detection, and system health monitoring, among other critical applications [12, 37]. In finance, for example, anomaly detection can signal fraudulent transactions or market manipulations, while in healthcare, it might indicate abnormal patient conditions that require immediate attention [10].
Initially, the field of time series anomaly detection heavily relied on statistical methods, grounded in robust mathematical frameworks, that have provided deep insights into data patterns and anomalies [49, 67]. These methods, while effective for simpler datasets, often assume specific data distributions or employ distance-based metrics to identify outliers. However, their application to the high-dimensional, complex nature of multivariate time series data reveals significant limitations. The primary challenge lies not merely in capturing the correlations within the data, but in effectively managing the intricacies associated with multi-correlated variables. These complexities can lead to an increased rate of false positives and missed detections, as traditional methods struggle to adequately interpret the nuanced interactions and dependencies that characterize modern datasets [23, 56].
The emergence of deep learning models, and their ability to learn high-level representations from massive amounts of data, have shown remarkable success in detecting anomalies in time series data [52]. These models excel in identifying deviations from frequently occurring patterns, significantly reducing false positives and improving detection accuracy [35, 39, 5]. A deeper exploration of machine learning models, including specific architectures and their successes across various applications, would provide valuable insights [36].
Among the deep learning architectures that have gained prominence, the transformer architecture introduced by [63] has revolutionized how models process sequential data. Unlike its predecessors, such as RNNs and LSTMs, the transformer leverages a self-attention mechanism, enabling it to capture long-range dependencies and temporal patterns in data without being hindered by the sequential processing bottleneck [38, 58, 30, 57, 32]. Its application in time series anomaly detection is particularly promising, offering a new perspective on modeling temporal anomalies. However, this area remains relatively unexplored, with significant potential for optimization and adaptation to the unique characteristics of time series data across different domains [60].
The Neural Architecture Search (NAS) has emerged as a transformative approach in automating the design of neural network architectures [71, 47, 13]. By systematically exploring a vast space of architectural configurations, NAS aims to identify models that achieve optimal performance for a given task. This automation is crucial in deep learning, where the design and tuning of models are both resource-intensive and require specialized knowledge [36, 18, 16]. Recent advancements in NAS have focused on improving its efficiency, reducing the computational resources required for the search process, and making state-of-the-art model architectures accessible for a broader range of applications [43, 71].
Incorporating multi-objective optimization into NAS, using algorithms like NSGA-II, represents a significant leap forward [21, 16]. For time series anomaly detection, where models must process large volumes of data accurately and with computational efficiency, the ability to find a nuanced balance between these competing objectives is crucial [66, 42, 71]. Despite the clear advantages, the application of multi-objective optimization in NAS for anomaly detection in time series is ripe for exploration and offers vast potential for groundbreaking research [36].
Our work seeks to harness the strengths of Neural Architecture Search (NAS) and transformer models to directly address the unique challenges associated with temporal data, crafting models that are finely tuned for anomaly detection in time series. This integration aims to develop architectures that excel in both performance and efficiency, leveraging the latest advancements in the field [15]. Given the nascent stage of combining NAS with transformer models for this purpose, our research not only contributes to the existing body of knowledge but also opens avenues for future exploration and innovation in anomaly detection methodologies [69]. By situating our study within the context of ongoing research efforts and emerging trends, we aim to advance the state of the art in anomaly detection within time series data and contribute meaningfully to the broader discourse in this rapidly evolving field.
@@ 3 Methodology
In this section, we delineate the comprehensive methodology employed in our study, designed to effectively detect and analyze anomalies in multivariate time series data using the TransNAS-TSAD framework. We commence with a clear definition of the problem, setting the stage for the subsequent methodological steps. This is followed by an in-depth discussion on data refinement techniques, ensuring that the data are appropriately pre-processed for our analysis. Next, we present the intricacies of our transformer architecture, highlighting its design and capabilities. A critical component of our approach is the neural architecture search (NAS), which employs the NSGA-II algorithm for optimizing the transformer model. That description is supplemented by an overview of the evolutionary process that iteratively refines our model architecture. We then transition to discussing advanced anomaly detection techniques, integrating adversarial elements into our model. The section culminates with a description of how we harness the full potential of TransNAS-TSAD for practical anomaly detection applications, illustrating the real-world implications of our research.
@@ 3.1 Problem Definition
Given a sequence of multivariate time series data $S=\{x_{1},x_{2},\ldots,x_{T}\}$, where each point $x_{t}$ is timestamped $t$ and resides in an $m$-dimensional space ($x_{t}\in\mathbb{R}^{m}$), our aims are to perform:
@@
Multi-Objective Transformer Model Optimization
At the heart of our methodology is the use of the NSGA-II based neural architecture search (NAS) for optimizing the transformer architecture, focusing on:
  • Maximizing Anomaly Detection Accuracy: We aim to enhance the classification performance (F1 score) for precise anomaly detection.
@@
Data Processing and Model Training
Prior to model training and classification, the time series data undergo a preprocessing phase where normalization is applied to ensure uniformity of scale across all series components. The normalization process is as follows:
$x_{t}\leftarrow\frac{x_{t}-\min(S)}{\max(S)-\min(S)+\epsilon},$
where $\min(S)$ and $\max(S)$ denote the component-wise minimum and maximum values observed in the training and testing time series data $S$, respectively. The term $\epsilon$ is a small constant introduced to prevent division by zero, ensuring numerical stability.
@@
Anomaly Scoring and Thresholding
Our model is designed to assign an anomaly score to each context window $W_{t}$, as opposed to direct binary labels. This score is derived from the reconstruction error of the window, where $O_{t}$ denotes the reconstructed output corresponding to $W_{t}$. A dynamic threshold $D$ is established based on the distribution of the scores computed on the training data, facilitating the distinction between normal and anomalous windows. The precise mechanism for determining the dynamic threshold $D$ and how it relates to the detection accuracy are described in subsequent sections.
@@ 3.2 TransNAS-TSAD Transformer Architecture
In the TransNAS-TSAD framework, the transformer architecture is specially adapted for time series anomaly detection. This involves a unique configuration of the standard transformer model to align with the specific characteristics of the time series data.
Data Augmentation Mechanisms Within the Model:
TransNAS-TSAD integrates specific data augmentation techniques within the model architecture. These techniques, which can be fine-tuned during training, include:
@@
Encoder Operations
The encoder in TransNAS-TSAD is designed with the following operations:
@@
    $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{\text{model}}}}\right)V.$ (1)
@@
    $\text{MultiHeadAtt}(Q,K,V)=\text{Concat}(H_{1},H_{2},\ldots,H_{h}),$ (2)
@@
Decoder Operations
The decoder mirrors the encoder’s complexity and includes additional transformations that integrate encoder memory, apply multi-head self-attention, and utilize feedforward neural networks, culminating in a final layer for accurate reconstruction. TransNAS-TSAD can utilize either a dual sequential decoder setup or an iterative approach, depending on the optimized configuration for maximum anomaly detection accuracy.